mes as old as he daughter. If the sum of their ages is thirty nine, find the aA mother is 3 years less than five tige of each.
Answer by Cromlix(4381)

Hi there,
A mother is 3 years less than
five times as old as he daughter.
M = 5D - 3
the sum of their ages is thirty nine.
M + D = 39
Replace M with 5D - 3 from top equation.
5D - 3 + D = 39
Collect like terms.
5D + D = 39 + 3
6D = 42
D = 7
Daughter is 7 years old
Mother is 32 years old
Hope this helps :-)